Q2
(a) . Given that in an economy
C = 0.8 (1-t) Y
t = 0.25
I = 900 - 50i
1
Expert's answer
2021-07-22T08:26:35-0400

The equilibrium level in the economy is derived from the IS And LM curve equilibrium position.

"Y=C+I+G"

"Y=0.8(1-0.25)Y+900-50i+800"

"Y=0.6Y+1700-50i"

"Y=0.25(1700-50i)"

"Y=4250-125i" (for IS curve equilibrium)


For LM curve equilibrium ;

"0.25Y-62.5i=500"

"Y=2000+250i"